Suppose you deposit $2,679.00 into an account today. In 12.00 years the account is worth $3,670.00. The account earned ____% per year.
Answer #1
Amount= Deposit * (1+rate)^Time
3670= 2679*(1+rate)^12
3670/2679= (1+rate)^12
1.369914147 =(1+rate)^12
1.369914^(1/12) =(1+rate)
1.026576004 =(1+rate)
Rate 0.026576
Rate 2.658% per year